# Fact-Checking at Scale: Multimodal AI for Authenticity and Context Verification in Online Media

### v1

@@ 2.1. Data Preprocessing
  • Media Encoding: All processed media content undergoes Base64 encoding to enable standardized API transmission across verification services. The system implements separate encoding functions for static images and video frames, ensuring consistent format handling regardless of input media type.
  • Integration and Standardization: The preprocessing module integrates metadata and processed media into a unified data structure compatible with downstream verification services. This includes organizing media content, extracting relevant textual features, and preparing data payloads for geolocation analysis, temporal verification, and content classification services. The standardized output format enables seamless integration with the evidence aggregation pipeline while maintaining data integrity throughout the verification process.
@@ 2.3.2. Where? (Location)
Figure 2. The pipeline of finding where images/videos are taken
@@ 2.3.3. When? (Time)
Figure 3. The pipeline of finding when images/videos are taken
@@ 3. Experimental Result
3.1. Implementation
The experiments were conducted on a computer system with the following specifications:
  • CPU: Xeon® Gold 6248R.
  • GPU: NVIDIA 4060 OC Edition 8GB GDDR6
  • RAM: 32GB.
The implementation was developed using Python programming language and the PyTorch deep learning framework (version 2.2.1) with CUDA version 11.8. The CPU was used for pre-processing tasks, model training, and evaluation, while the GPU was employed to accelerate computationally intensive deep learning algorithms.
@@ 3.4.1. Main-task
Team TOTAL
Team 1 844.22
Team 2 487.19
Team 3 295.86
Ours 644.65
@@ Acknowledgements.
All intellectual property rights arising from the work shall belong solely and exclusively to the National Institute of Information and Communications Technology (NICT), and may be used by NICT for any purpose, including but not limited to commercial applications, without any claim, restriction, or obligation from or to non-NICT contributors.

### v2

@@ 2.1. Data Preprocessing
  • Media Encoding: All processed media content undergoes Base64 encoding to enable standardized API transmission across verification services. The system implements separate encoding functions for static images and video frames, ensuring consistent format handling regardless of input type.
  • Integration and Standardization: The preprocessing module integrates metadata and processed media into a unified data structure compatible with downstream verification services. This includes organizing media content, extracting relevant textual features, and preparing data payloads for geolocation analysis, temporal verification, and content classification services. The standardized output format enables seamless integration with the evidence aggregation pipeline while maintaining data integrity throughout the verification.
@@ 2.3.2. Where? (Location)
Figure 2. The pipeline of finding (A) where and (B) when images/ videos are taken
@@ 3.3.1. Main-task
Team 1 Team 2 Team 3 Ours
Score 844.22 487.19 295.86 644.65
@@ Acknowledgements.
The authors gratefully acknowledge the financial support provided by FPT Software for covering the research registration fee associated with this paper. All intellectual property rights arising from the work shall belong solely and exclusively to the National Institute of Information and Communications Technology (NICT), and may be used by NICT for any purpose, including but not limited to commercial applications, without any claim, restriction, or obligation from or to non-NICT contributors.